OpenStudy (anonymous):

An skater is able to spin at 4 rev/s. What is her angular velocity in rad/s? What is the centripetal acceleration of her arms if they are 0.85 meters from the axis of rotation?

OpenStudy (michele_laino):

here we have to apply this formula: \[\omega = 2\pi f \cong 6.28 \cdot 4 = ...?\] being \(f=4 \,s^{-1}\)

OpenStudy (anonymous):

Where does the 6.28 came from?

OpenStudy (michele_laino):

I have used an approximated result, here is my step: \[2\pi = 2 \cdot 3.14159... \cong 2 \cdot 3.14 = 6.28\]

OpenStudy (anonymous):

Ok. thanks!